OpenStudy (anonymous):

the width "w" of a rectangular pool is x+4. the are of the pool is 2x^3-29x+12. what is an expression for the length of the pool?

OpenStudy (anonymous):

weird way to ask you to factor \[2x^2-29x+12=(x+4)(\text{something})\]

OpenStudy (anonymous):

@satellite73 can you come moderate a discussion, please? I'm arguing with someone and I don't know why.

OpenStudy (anonymous):

i am hoping there is a typo here

OpenStudy (campbell_st):

try a little polynomial division

OpenStudy (anonymous):

@Numb3r1 where?

OpenStudy (anonymous):

campbell_st how do i do it?

OpenStudy (anonymous):

ooh it is cubed!!!

OpenStudy (anonymous):

http://openstudy.com/study#/updates/518eeaf0e4b0cf6dd8f45beb There have also been a couple of private messages, but the gist of it is there.

OpenStudy (anonymous):

\[2x^3-29x+12=(x+4)(\text{something})\]

OpenStudy (anonymous):

you can figure it out either by dividing or by thinking

OpenStudy (anonymous):

Divide as if you were doing long division (multiply to fit the first value, then subtract, then do the same for the next values, etc.)

OpenStudy (anonymous):



OpenStudy (campbell_st):

here is the division  hope this helps